1.ariel has $5.50 in dimes and quaters. she has 8 more quaters than dimes. How many quaters does she have?
2. tenisha and dameisha have the same number of coins. Tenisha has only dimes and dameisha has only quaters. if damiesha has $1.80 more than tenisha does, how much does each have?

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8405)

Q=D+8 8 MORE QUARTERS THAN DIMES
NOW ADD THE VALUESOF THE QUARTERS & DIMES WE GET
25(D+8)+10D=550
25D+200+10D=550
35D=550-200
35D=350
D=10 SOLUTION FOR THE NUMBER OF DIMES.
THEN THERE ARE Q=10+8
Q=18 SOLUTION FORTHE NUMBER OF QUARTERS.
PROOF
550=18*25+10*10
550=450+100
550=550

Answer by stanbon(29602)

1.ariel has $5.50 in dimes and quarters. she has 8 more quarters than dimes. How many quaters does she have?
Let # of dimes be x : value of these is 10x cents
Number of quarters is "x+8"; value of these is 25(x+8)=25x+200 cents
EQUATION:
value + value = 550 cents
10x+25x+200=550
35x=350
x=10 (number of dimes)
x+8=18 (number of quarters)
-----------------
2. tenisha and dameisha have the same number of coins. Tenisha has only dimes and dameisha has only quaters. if damiesha has $1.80 more than tenisha does, how much does each have?
Tenisha DATA:
# of dimes is "x"; value of these is 10x cents
Dameisha DATA:
# of quarters is "x"; value of these is 25x cents
EQUATION:
Tenisha value + 180 = damiesha value
10x+180=25x
15x=180
x=12
Tenisha has 12*10=$1.20
Dameisha has 12*25=$3.00
---------------
Cheers,
Stan H.